# Stochastic Particle Acceleration during Pressure-Anisotropy-Driven Magnetogenesis in the Pre-Structure Universe

### 05 Mar 2026 v1

@@ Abstract
We investigate whether stochastic acceleration associated with pressure-anisotropy-driven magnetogenesis can generate a dynamically significant population of cosmic rays (CRs) prior to nonlinear structure formation. As magnetic fields amplify in the early Universe, the associated increase in gyrofrequency enhances pitch-angle scattering, potentially shortening the stochastic acceleration time. We derive an analytic criterion for efficient cosmological acceleration by comparing the acceleration timescale with the Hubble time, which defines a critical magnetic field and a corresponding CR turn-on redshift $z_{\rm on}$. For representative parameters, we find $z_{\rm on}\sim 1.7$. To quantify the resulting particle population, we solve a Fokker-Planck equation for the isotropic proton distribution in the redshift interval $z=10\rightarrow z_{\rm on}$. Throughout most of this epoch, adiabatic expansion dominates over stochastic energization and the distribution remains close to a cooling Maxwellian. However, as the system approaches the turn-on epoch, the stochastic acceleration time decreases, allowing a mild suprathermal tail to develop. Even under optimistic assumptions corresponding to the strong-scattering limit, the maximum attainable proton energy reaches at most $\mathcal{O}(10^{2})\,\mathrm{GeV}$. These results indicate that efficient CR production in the intergalactic medium is intrinsically tied to the onset of structure-formation shocks, while earlier microinstability-driven stochastic processes can provide at most a modest pre-acceleration.
I Introduction
@@
Cosmic rays (CRs) are a ubiquitous nonthermal component of the Universe, observed across environments ranging from galaxies to galaxy clusters. In the Milky Way, the CR energy density is of order $\sim 1~{\rm eV\,cm^{-3}}$, comparable to that of the thermal interstellar medium, implying that CRs play a dynamically significant role in galactic systems [e.g., 21]. It is widely accepted that supernova remnants accelerate CRs up to $\sim 10^{15}$ eV via diffusive shock acceleration (DSA) [e.g., 2, 3, 5], and multiwavelength observations provide strong evidence for relativistic particle acceleration at such shocks [e.g., 16, 22, 1]. In some cases, CR precursors predicted by DSA have been observationally identified [24]. In the context of large-scale structure formation, DSA at collisionless shocks is likewise regarded as a primary mechanism responsible for accelerating particles to relativistic energies. Structure-formation shocks with velocities $u_{s}\sim 10^{2}$$10^{3}~{\rm km\,s^{-1}}$ naturally arise during hierarchical clustering, and numerical simulations support their role in generating CR populations in the intracluster and intergalactic media [e.g., 20, 28, 26, 14, 33, 34, 15, 29, 9, 10, 11].
@@
However, an important open question concerns the origin of the first nonthermal particles. Cosmological simulations indicate that a large fraction of gas thermalization and cosmic-ray production occurs at structure-formation shocks emerging at relatively low redshifts ($z\lesssim{\rm a~few}$) [e.g., 28]. If CR production is tied exclusively to such shocks, then a pre-existing suprathermal seed population may not be required. Nevertheless, energetic events at earlier epochs may still generate the first CRs. For instance, early shock-driven scenarios associated with the first supernova explosions at $z\sim 20$ have been proposed as possible sources [e.g., 19, 23]. Another possibility arises during the subsequent pre-structure epoch, before large-scale structure formation enters the nonlinear regime and cosmological shocks become widespread ($z\lesssim\text{a few}$). During this period, the Universe undergoes a prolonged phase of weak magnetization, and magnetic fields are expected to grow gradually through various amplification processes [e.g., 27, 35]. If weak but cumulative acceleration processes operate during this intermediate redshift interval ($z\sim 10\rightarrow{\rm a~few}$), they could establish a low-level CR background that modifies subsequent injection and acceleration at structure-formation shocks. Determining whether such a pre-shock population exists is therefore relevant for understanding both the efficiency of DSA and the thermal history of the early intergalactic medium.
@@
In parallel, recent studies of cosmological magnetogenesis have emphasized the role of pressure-anisotropy-driven plasma instabilities in amplifying magnetic fields in weakly magnetized environments [e.g., 30, 31, 32, 6, 12]. As the magnetic field grows, the associated proton gyrofrequency increases, enhancing pitch-angle scattering and potentially reducing the characteristic acceleration time of stochastic (second-order Fermi) processes operating in turbulence [e.g., 7, 4, 25]. This raises a natural question: can instability-assisted stochastic acceleration during cosmological magnetogenesis generate a dynamically significant CR population prior to the emergence of structure-formation shocks?
@@
In this work, we investigate this possibility by combining (i) an analytic acceleration-time criterion based on the comparison between the stochastic acceleration time and the Hubble time, and (ii) a Fokker-Planck model for the isotropic proton distribution. This approach allows us to quantify both the redshift at which stochastic acceleration becomes cosmologically viable and the magnitude of any resulting nonthermal population. The central question we address is whether instability-enhanced scattering during cosmological magnetogenesis can reduce the stochastic acceleration time below the cosmological expansion timescale over any redshift interval prior to the widespread emergence of structure-formation shocks. In particular, we derive a critical condition for cosmologically viable second-order Fermi acceleration and follow the redshift evolution of the proton distribution under the combined effects of stochastic diffusion and adiabatic expansion. This framework enables a self-consistent assessment of (i) the epoch at which stochastic acceleration may become competitive with cosmic expansion, and (ii) the magnitude of any resulting suprathermal population.
@@
The paper is organized as follows. In Section II, we derive the stochastic acceleration timescale in the instability-mediated magnetogenesis framework and obtain an analytic criterion for the turn-on of cosmological acceleration. In Section III, we solve a minimal Fokker-Planck equation to quantify the pre-structure stochastic heating of the proton distribution. In Section IV, we compare the stochastic acceleration timescale with DSA in the Bohm-diffusion limit. Section V summarizes our results and discusses their cosmological implications.
@@
II Stochastic Acceleration During Pressure-Anisotropy-Driven Magnetogenesis
Magnetic-field amplification mediated by pressure-anisotropy-driven instabilities can substantially modify particle transport properties in weakly magnetized plasmas. Since magnetic fields play a central role in modulating particle scattering, the growth of cosmic magnetic fields prior to the onset of large-scale structure formation may naturally provide conditions for stochastic particle acceleration. In this section, we examine whether particles can be accelerated on cosmological timescales through second-order Fermi processes operating in turbulence generated during magnetogenesis.
II.1 Acceleration Timescale
@@
We adopt a stochastic acceleration scenario in which particles gain energy through repeated scattering off moving magnetic irregularities. The characteristic acceleration timescale can be written as
$t_{\rm acc}(E,z)\approx\left(\frac{c}{v_{\rm tur}(z)}\right)^{2}\frac{\lambda(E,z)}{c},$ (1)
where $v_{\rm tur}(z)$ is the turbulent velocity at the outer scale, $\lambda(E,z)$ is the particle mean free path, and $c$ is the speed of light. For relativistic particles, the mean free path is determined by the effective scattering frequency,
$\lambda(E,z)\approx\frac{v(E)}{\nu_{\rm eff}(z)}.$ (2)
@@
$\nu_{\rm eff}(z)\equiv\nu_{ii}(z)+\nu_{\rm scatt}(z),$ (3)
where $\nu_{ii}$ is the Coulomb collision frequency and $\nu_{\rm scatt}$ represents the scattering induced by pressure-anisotropy-driven instabilities.
Following the instability-mediated magnetogenesis framework [e.g., 12], the scattering frequency in the regime where plasma instabilities dominate can be expressed as
$\nu_{\rm scatt}\sim\left(|\Delta|-2\beta^{-1}\right)^{3/2}\omega_{\rm cp},$ (4)
where $\Delta$ is the pressure anisotropy factor, $\beta$ is the plasma beta parameter, and $\omega_{\rm cp}=eB/m_{p}c$ is the proton gyrofrequency associated with the evolving magnetic field. Substituting Eq. (2) into Eq. (1), the acceleration timescale becomes
$t_{\rm acc}(z)\approx\left(\frac{c}{v_{\rm tur}(z)}\right)^{2}\frac{v(E)}{c\nu_{\rm eff}(z)}.$ (5)
This expression highlights that magnetic-field amplification, through its impact on $\omega_{\rm cp}$ and $\nu_{\rm scatt}$, can significantly shorten the acceleration time. Throughout this work, we adopt the scattering-dominated regime, consistent with the transition discussed in previous studies [6, 12]. In the very early stages of magnetic-field amplification, collisional effects can in principle influence the plasma dynamics. However, as the magnetic field grows and the proton gyrofrequency $\omega_{\rm cp}\propto B$ increases, the ratio $\nu_{\rm coll}/\omega_{\rm cp}\propto(1+z)^{3}B^{-1}$ rapidly decreases. For redshifts $z\lesssim 10$, where the baryon density has already declined substantially and magnetic amplification has progressed, the condition $\nu_{\rm coll}/\omega_{\rm cp}\ll 1$ is naturally satisfied. In this regime, pitch-angle scattering mediated by plasma instabilities dominates over Coulomb collisions ($\nu_{ii}\ll\nu_{\rm scatt}$). The effective scattering rate therefore satisfies $\nu_{\rm eff}\approx\nu_{\rm scatt}$, making the stochastic acceleration framework adopted here self-consistent.
@@
$H(z)=H_{0}\sqrt{\Omega_{m}(1+z)^{3}+\Omega_{\Lambda}}.$ (7)
@@
$\frac{t_{\rm acc}}{t_{H}}\approx\left(\frac{c}{v_{\rm tur}}\right)^{2}\left(\frac{v(E)}{c}\right)\frac{H(z)}{\nu_{\rm eff}(z)}.$ (9)
When this ratio falls below unity, particles can be accelerated to relativistic energies within a cosmological expansion time. Conversely, if $t_{\rm acc}>t_{H}$, the expansion of the universe limits the achievable particle energies.
II.2 Implications for Pre-Shock CR Production
Eqs. (1)-(9) suggest that instability-driven scattering during magnetogenesis may create a channel for CR production prior to structure formation shocks. As the magnetic field grows, the associated increase in gyrofrequency enhances the scattering rate, thereby reducing the acceleration timescale. This feedback implies that even moderately amplified magnetic fields could enable stochastic acceleration in the early universe.
In this framework, the key quantity controlling the onset of acceleration is the ratio $H(z)/\nu_{\rm eff}(z)$. If instability-mediated scattering becomes sufficiently strong, stochastic acceleration may operate well before the emergence of collisionless shocks, potentially establishing a seed population of CRs that later participate in DSA. The results presented here therefore motivate a quantitative exploration of particle acceleration during cosmological magnetogenesis and its implications for the origin of the earliest nonthermal particle populations.
To evaluate the viability of stochastic acceleration during pressure-anisotropy-driven magnetogenesis, we adopt a minimal self-consistent closure of the background plasma parameters. Throughout this section we restrict our analysis to protons; extension to electrons will be discussed in future work.
Turbulent velocity.

### 01 Apr 2026 v2

@@ Abstract
We investigate whether stochastic acceleration associated with pressure-anisotropy-driven magnetogenesis can generate a dynamically significant population of cosmic rays (CRs) prior to nonlinear structure formation. As magnetic fields amplify in the early Universe, the associated increase in gyrofrequency enhances pitch-angle scattering, potentially shortening the stochastic acceleration time. We derive an analytic criterion for efficient cosmological acceleration by comparing the acceleration timescale with the Hubble time, which defines a critical magnetic field and a corresponding CR turn-on redshift $z_{\rm on}$. For representative parameters, we find $z_{\rm on}\sim 1.7$. To quantify the resulting particle population, we solve a Fokker-Planck equation for the isotropic ion (proton) distribution in the redshift interval $z=10\rightarrow z_{\rm on}$, including Coulomb energy losses in a fully ionized intergalactic medium. Throughout most of this epoch, adiabatic expansion dominates over stochastic energization, and Coulomb cooling efficiently thermalizes low-energy particles, introducing an effective injection threshold at energies of order ${\mathcal{O}}(10)$ keV. As a result, the distribution remains close to a cooling Maxwellian, and the formation of a suprathermal tail is strongly suppressed even in the presence of a pre-existing nonthermal component. Even under optimistic assumptions corresponding to the strong-scattering limit, the maximum attainable ion energy reaches at most $\mathcal{O}(10^{2})$ GeV. These results indicate that efficient CR production in the intergalactic medium is intrinsically tied to the onset of structure-formation shocks, while earlier microinstability-driven stochastic processes can provide at most a modest pre-acceleration.
1 Introduction
@@
Cosmic rays (CRs) are ubiquitous nonthermal components of the Universe, observed across environments ranging from galaxies to galaxy clusters. In the Milky Way, the CR energy density is of order $\sim 1~{\rm eV\,cm^{-3}}$, comparable to that of the thermal interstellar medium, implying that CRs play a dynamically significant role in galactic systems [e.g., 22]. It is widely accepted that supernova remnants accelerate CRs up to $\sim 10^{15}$ eV via diffusive shock acceleration (DSA) [e.g., 2, 3, 5], and multiwavelength observations provide strong evidence for relativistic particle acceleration at such shocks [e.g., 17, 23, 1]. In some cases, CR precursors predicted by DSA have been observationally identified [25]. In the context of large-scale structure formation, DSA at collisionless shocks is likewise regarded as a primary mechanism responsible for accelerating particles to relativistic energies. Structure-formation shocks with velocities $u_{s}\sim 10^{2}$$10^{3}~{\rm km\,s^{-1}}$ naturally arise during hierarchical clustering, and numerical simulations support their role in generating CR populations in the intracluster and intergalactic media [e.g., 21, 29, 27, 15, 36, 37, 16, 31, 10, 11, 12].
@@
However, an important open question concerns the origin of the first nonthermal particles. Cosmological simulations indicate that a large fraction of gas thermalization and cosmic-ray production occurs at structure-formation shocks emerging at relatively low redshifts ($z\lesssim{\rm a~few}$) [e.g., 29]. If CR production is tied exclusively to such shocks, then a pre-existing suprathermal seed population may not be required. Nevertheless, energetic events at earlier epochs may still generate the first CRs. For instance, early shock-driven scenarios associated with the first supernova explosions at $z\sim 20$ have been proposed as possible sources [e.g., 20, 24]. Another possibility arises during the subsequent pre-structure epoch, before large-scale structure formation enters the nonlinear regime and cosmological shocks become widespread ($z\lesssim\text{a few}$). During this period, the Universe undergoes a prolonged phase of weak magnetization, and magnetic fields are expected to grow gradually through various amplification processes [e.g., 28, 38]. If weak but cumulative acceleration processes operate during this intermediate redshift interval ($z\sim 10\rightarrow{\rm a~few}$), they could establish a low-level CR background that modifies subsequent injection and acceleration at structure-formation shocks. Determining whether such a pre-shock population exists is therefore relevant for understanding both the efficiency of DSA and the thermal history of the early intergalactic medium.
@@
In parallel, recent studies of cosmological magnetogenesis have emphasized the role of pressure-anisotropy-driven plasma instabilities in amplifying magnetic fields in weakly magnetized environments [e.g., 32, 33, 34, 6, 13]. As the magnetic field grows, the associated ion (proton) gyrofrequency increases, enhancing pitch-angle scattering and potentially reducing the characteristic acceleration time of stochastic (second-order Fermi) processes operating in turbulence [e.g., 7, 4, 26]. This raises a natural question: can instability-assisted stochastic acceleration during cosmological magnetogenesis generate a dynamically significant CR population prior to the emergence of structure-formation shocks?
@@
In this work, we investigate this possibility by combining (i) an analytic acceleration-time criterion based on the comparison between the stochastic acceleration time and the Hubble time, and (ii) a Fokker-Planck model for the isotropic ion distribution. This approach allows us to quantify both the redshift at which stochastic acceleration becomes cosmologically viable and the magnitude of any resulting nonthermal population. The central question we address is whether instability-enhanced scattering during cosmological magnetogenesis can reduce the stochastic acceleration time below the cosmological expansion timescale over any redshift interval prior to the widespread emergence of structure-formation shocks. In particular, we derive a critical condition for cosmologically viable second-order Fermi acceleration and follow the redshift evolution of the ion distribution under the combined effects of stochastic diffusion and adiabatic expansion. This framework enables a self-consistent assessment of (i) the epoch at which stochastic acceleration may become competitive with cosmic expansion, and (ii) the magnitude of any resulting suprathermal population. Energy losses due to Coulomb interactions are known to play an important role in shaping the low-energy spectrum of charged particles in diffuse astrophysical plasmas [e.g., 9, 30, 26]. In particular, Coulomb collisions can efficiently thermalize sub-relativistic particles and suppress the formation of nonthermal tails in weakly collisional environments. However, their impact on stochastic acceleration in the cosmological context—especially during the magnetogenesis phase prior to nonlinear structure formation—has not been systematically quantified. In this work, we incorporate Coulomb losses into both the analytic framework and the Fokker–Planck model, and examine how collisional cooling modifies particle injection and the resulting nonthermal population.
@@
The paper is organized as follows. In Section 2, we derive the stochastic acceleration timescale within the instability-mediated magnetogenesis framework and obtain an analytic criterion for the cosmological viability of stochastic acceleration by considering adiabatic losses associated with cosmic expansion. This provides an optimistic upper bound on the efficiency of particle acceleration and defines a critical magnetic field and a corresponding turn-on redshift. In Section 3, we extend this analysis by including Coulomb energy losses in a fully ionized intergalactic medium. In Section 4, we solve a Fokker–Planck equation to quantify the resulting ion distribution. This allows us to assess the impact of collisional cooling on particle injection and to identify an effective energy threshold for stochastic acceleration. In Section 5, we compare the stochastic acceleration timescale with that of DSA in the Bohm-diffusion limit. Finally, Section 6 summarizes our results and discusses their cosmological implications.
@@
2 Cosmological Viability of Stochastic Acceleration During Pressure-Anisotropy-Driven Magnetogenesis
Magnetic-field amplification mediated by pressure-anisotropy-driven instabilities can substantially modify particle transport properties in weakly magnetized plasmas. Since magnetic fields play a central role in regulating particle scattering, the growth of cosmic magnetic fields prior to the onset of large-scale structure formation may provide conditions for stochastic particle acceleration. In this section, we assess the cosmological viability of such acceleration by deriving a necessary condition set by the expansion of the universe. Specifically, we compare the stochastic acceleration timescale with the Hubble time, which defines the minimum requirement for particle energization on cosmological timescales. This approach provides an optimistic estimate of the acceleration efficiency and serves as a baseline against which additional physical effects, such as collisional energy losses, can be evaluated in subsequent sections. Throughout this work, we focus on ions, specifically protons, as the dominant baryonic component of the plasma, which sets the relevant mass and gyrofrequency scales for the acceleration process.
2.1 Acceleration Timescale
@@
We adopt a stochastic acceleration scenario in which particles gain energy through repeated scattering off moving magnetic irregularities. In general, the acceleration timescale depends on the velocity fluctuations at the scale with which particles interact, i.e., at scales comparable to the particle mean free path $\lambda(E,z)$. The acceleration timescale can therefore be expressed schematically as
$t_{\rm acc}(E,z)\sim\left(\frac{c}{v_{\lambda}(z)}\right)^{2}\frac{\lambda(E,z)}{c},$ (2.1)
where $v_{\lambda}(z)$ denotes the characteristic turbulent velocity at the scale $\lambda$. In the present work, we adopt an optimistic approximation in which particles are assumed to sample velocity fluctuations comparable to the outer-scale turbulent velocity, $v_{\lambda}\sim v_{\rm tur}$. Under this assumption, the acceleration timescale reduces to
$t_{\rm acc}(E,z)\sim\left(\frac{c}{v_{\rm tur}(z)}\right)^{2}\frac{\lambda(E,z)}{c},$ (2.2)
which should be regarded as a lower-bound estimate of the true acceleration time. In realistic turbulence, where $v_{\lambda}<v_{\rm tur}$ for $\lambda$ below the outer scale, the actual acceleration time is expected to be longer, thereby further suppressing particle acceleration.
The mean free path is determined by the effective scattering frequency,
$\lambda(E,z)\approx\frac{v(E)}{\nu_{\rm eff}(z)}.$ (2.3)
@@
$\nu_{\rm eff}(z)\equiv\nu_{ii}(z)+\nu_{\rm scatt}(z),$ (2.4)
where $\nu_{ii}$ is the ion-ion (proton-proton) Coulomb collision frequency and $\nu_{\rm scatt}$ represents the scattering induced by pressure-anisotropy-driven instabilities. Following the instability-mediated magnetogenesis framework [e.g., 13], the scattering frequency in the regime where plasma instabilities dominate can be expressed as
$\nu_{\rm scatt}\sim\left(|\Delta|-2\beta^{-1}\right)^{3/2}\omega_{\rm ci},$ (2.5)
where $\Delta$ is the pressure anisotropy factor, $\beta$ is the plasma beta parameter, and $\omega_{\rm ci}=eB/m_{i}c$ is the ion gyrofrequency associated with the evolving magnetic field. This expression characterizes the pitch-angle scattering rate induced by pressure-anisotropy-driven microinstabilities, and is most directly applicable to particles with velocities comparable to the thermal speed. For higher-energy particles, the effective scattering rate is expected to depend on particle rigidity, leading to an energy-dependent diffusion coefficient. In the present work, we adopt Eq. (2.5) as an effective scattering rate that sets the normalization of stochastic acceleration. This corresponds to an optimistic assumption that maximizes the scattering efficiency. Any additional energy dependence that reduces the scattering rate at high energies would increase the acceleration time, thereby further suppressing particle acceleration.
Substituting Eq. (2.3) into Eq. (2.2), the acceleration timescale becomes
$t_{\rm acc}(z)\approx\left(\frac{c}{v_{\rm tur}(z)}\right)^{2}\frac{v(E)}{c\nu_{\rm eff}(z)}.$ (2.6)
This expression highlights that magnetic-field amplification, through its impact on $\omega_{\rm ci}$ and $\nu_{\rm scatt}$, can significantly shorten the acceleration time. Throughout this work, we adopt the scattering-dominated regime, consistent with the transition discussed in previous studies [6, 13]. In the very early stages of magnetic-field amplification, collisional effects can in principle influence the plasma dynamics. However, as the magnetic field grows and the ion gyrofrequency $\omega_{\rm ci}\propto B$ increases, the ratio $\nu_{ii}/\omega_{\rm ci}\propto(1+z)^{3}B^{-1}$ rapidly decreases. For redshifts $z\lesssim 10$, where the baryon density has already declined substantially and magnetic amplification has progressed, the condition $\nu_{ii}/\omega_{\rm ci}\ll 1$ is naturally satisfied. In this regime, pitch-angle scattering mediated by plasma instabilities dominates over Coulomb collisions ($\nu_{ii}\ll\nu_{\rm scatt}$). The effective scattering rate therefore satisfies $\nu_{\rm eff}\approx\nu_{\rm scatt}$, making the stochastic acceleration framework adopted here self-consistent.
@@
$H(z)=H_{0}\sqrt{\Omega_{m}(1+z)^{3}+\Omega_{\Lambda}}.$ (2.8)
@@
$\frac{t_{\rm acc}}{t_{H}}\approx\left(\frac{c}{v_{\rm tur}}\right)^{2}\left(\frac{v(E)}{c}\right)\frac{H(z)}{\nu_{\rm eff}(z)}.$ (2.10)
When this ratio falls below unity, particles can be accelerated to relativistic energies within a cosmological expansion time. Conversely, if $t_{\rm acc}>t_{H}$, the expansion of the universe limits the achievable particle energies. From Eq. (2.10), one finds that the acceleration rate is related to the effective scattering rate as
$t_{\rm acc}^{-1}\sim\left(\frac{v_{\rm tur}}{c}\right)^{2}\left(\frac{c}{v}\right)\nu_{\rm eff},$ (2.11)
which reduces to $t_{\rm acc}^{-1}\sim(v_{\rm tur}/c)^{2}\nu_{\rm eff}$ in the relativistic limit ($v\sim c$). This shows that even in the presence of strong scattering, the acceleration efficiency is suppressed by the small ratio $(v_{\rm tur}/c)^{2}$. As a result, rapid pitch-angle scattering does not necessarily imply efficient particle acceleration.
Equation (2.9) provides a necessary condition for stochastic acceleration on cosmological timescales. However, it is not sufficient, as additional energy loss processes can impose more stringent constraints, particularly at low energies. In the low-energy regime, the characteristic timescales exhibit a clear hierarchy. Pitch-angle scattering is dominated by instability-driven processes, satisfying $\nu_{ii}\ll\nu_{\rm scatt}$, while Coulomb interactions can still strongly affect the energy evolution of particles. In particular, the Coulomb energy loss rate can be comparable to or exceed the stochastic acceleration rate, $t_{C}^{-1}\gtrsim t_{\rm acc}^{-1}$, thereby suppressing efficient energization of sub-relativistic particles. A more general requirement for particle acceleration is therefore
$t_{\rm acc}(E,z)<\min\left[t_{H}(z),t_{C}(E,z)\right],$ (2.12)
where $t_{C}\equiv E/|\dot{E}_{C}|$ is the characteristic Coulomb energy loss timescale. This condition highlights that, even in the presence of efficient pitch-angle scattering, stochastic acceleration can be strongly suppressed at low energies due to collisional cooling. In this regime, particles undergo frequent scattering but gain energy only slowly, while Coulomb losses efficiently remove energy, preventing their injection into the nonthermal population. In the present analysis, we first use the comparison with the Hubble time to identify the cosmological regime where acceleration may become viable. The impact of Coulomb losses is incorporated in Section 3, where the acceleration condition is refined to include energy-dependent cooling effects that introduce an effective injection threshold for stochastic acceleration.
2.2 Implications for Pre-Shock CR Production
Eqs. (2.2)-(2.10) suggest that instability-driven scattering during magnetogenesis may create a channel for CR production prior to structure formation shocks. As the magnetic field grows, the associated increase in gyrofrequency enhances the scattering rate, thereby reducing the acceleration timescale. This feedback implies that even moderately amplified magnetic fields could enable stochastic acceleration in the early universe.
In this framework, the key quantity controlling the onset of acceleration is the ratio $H(z)/\nu_{\rm eff}(z)$. If instability-mediated scattering becomes sufficiently strong, stochastic acceleration may operate well before the emergence of collisionless shocks, potentially establishing a seed population of CRs that later participate in DSA. The results presented here therefore motivate a quantitative exploration of particle acceleration during cosmological magnetogenesis and its implications for the origin of the earliest nonthermal particle populations. To evaluate the viability of stochastic acceleration during pressure-anisotropy-driven magnetogenesis, we adopt a minimal self-consistent closure of the background plasma parameters.